OpenStudy (anonymous):

Electrical energy (E) is measured in kilowatt hours (kwh) using the formula E = P · t, where P is the power in kilowatts and t is the time in hours. If 128 kwh of electrical energy are created over a period of 16 hours, how many kilowatts of power are used each hour?

OpenStudy (anonymous):

@Cydney_Morgan @Destinymasha @aaronq @johnweldon1993 @undeadknight26 @Whitemonsterbunny17 @ShayleeCutie @Ashleyisakitty

OpenStudy (kirbykirby):

\(E=P\cdot t\) \(128 ~kwh = P(16~h)\\ P=\dfrac{128~kwh}{16~h}=?~kw\)

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

yup :)